OpenStudy (anonymous):

whats the simplified version of x^2 over x^6?

OpenStudy (anonymous):

x^-4 = 1/x^4

OpenStudy (anonymous):

if you can't see it right away just simplify it

OpenStudy (anonymous):

x^2= x*x

OpenStudy (anonymous):

x^6 = x*x*x*x*x*x

OpenStudy (anonymous):

anything over itself is one

OpenStudy (anonymous):

besides zero

OpenStudy (anonymous):

so whats the answer?

OpenStudy (anonymous):

google it